# Michael–Simon inequality for anisotropic energies close to the area via multilinear Kakeya-type bounds

### v1

@@ Abstract.
While a monotonicity formula is not available for critical points [4], when $k=2$ and $n=3$ we show that the Michael–Simon inequality holds if $F$ is convex and close to $1$ (in $C^{1}$), meaning that $\mathcal{F}$ is close to the usual area.
@@ 1.1. Setting and main result
In particular, Eq. 1 holds for a possibly different constant $C(F)$.
@@ 1.2. New variants of multilinear Kakeya in dimension two
Given $S,T\in W^{1,1}(\mathbb{R}^{2},\mathbb{R}^{2})$, assume that $S^{x},T^{y}\geq 0$ and $\det(S,T)\geq 0$ a.e. Then
@@
Also, let $\chi:\mathbb{R}^{2}\to[0,1]$ be a Borel function supported in a bounded set. Assuming $S^{x},T^{y}\geq 0$ then, for some universal constant $C>0$, we have
(4) $\begin{aligned} \int_{\mathbb{R}^{2}}\chi\min\{S^{P},T^{P}\}&\leq C\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}(|S|+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(|T|+\lvert\operatorname{div}T\rvert)\Big]^{1/2}\\ &\quad+\int_{\mathbb{R}^{2}}(C\lvert\operatorname{div}S\rvert+C\lvert\operatorname{div}T\rvert+S^{N}+T^{N}).\end{aligned}$
The same holds if $S^{x},S^{y},T^{x},T^{y},\operatorname{div}S,\operatorname{div}T$ are just real-valued measures on the plane with finite total variation (provided that $S^{x},T^{y}\geq 0$).
@@ 1.3. General results in arbitrary dimension
While we study autonomous integrands $F(P)$, for simplicity and in order to have scale-invariant bounds when possible, in general one can consider non-autonomous ones, of the form $F(x,P)$. This generalization is necessary in order to look at anisotropic integrands on manifolds $M^{n}$, where $F$ is defined on the Grassmannian bundle of $k$-planes and thus locally takes this form. Compared to the autonomous case, the general case just introduces an error term which is easily handled (and can be locally incorporated in a reduced form of $\delta^{F}V$, namely the right-hand side of Eq. 6). Straightforward modifications show that all the stated consequences of Michael–Simon still hold true, assuming this bound in a local form such as
@@ Acknowledgements
As already mentioned, this paper uses many ideas due to Frederick J. Almgren Jr. The authors are very grateful to Jean Taylor, who shared with them the unpublished manuscript of Almgren. A.P.’s research is funded by the European Research Council (ERC) through StG 101165368 “MAGNETIC”. G.D.P.’s research is funded by the European Research Council (ERC) through CoG 101169953 “RISE”. [^1] Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council.
@@
Up to a subsequence, we can assume that $|\delta^{F}V_{i}|\rightharpoonup\nu$ for a suitable Radon measure $\nu$. For $|V|$-a.e. $x$, the rectifiable varifold $V$ admits an approximate tangent plane and there exists $C>0$, depending on $x$, such that [^2] Recall that, for any two Radon measures $\nu$ and $\mu$, we have $\limsup_{r\to 0}\frac{\nu(B_{r}(x))}{\mu(B_{r}(x))}<+\infty\quad\text{for $\mu$-a.e.\ $x$.}$
@@ 3. Nonlinear inequalities for vector fields on the plane
Also, let $\chi:\mathbb{R}^{2}\to[0,1]$ be a Borel function supported in a bounded set. Assuming $S^{x},T^{y}\geq 0$ then, for some universal constant $C>0$, we have
(12) $\begin{aligned} \int_{\mathbb{R}^{2}}\chi\min\{S^{P},T^{P}\}&\leq C\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}(|S|+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(|T|+\lvert\operatorname{div}T\rvert)\Big]^{1/2}\\ &\quad+\int_{\mathbb{R}^{2}}(C\lvert\operatorname{div}S\rvert+C\lvert\operatorname{div}T\rvert+S^{N}+T^{N}).\end{aligned}$
The same holds if $S^{x},S^{y},T^{x},T^{y},\operatorname{div}S,\operatorname{div}T$ are just real-valued measures on the plane with finite total variation (provided that $S^{x},T^{y}\geq 0$).
@@
In the general case of measures, quantities such as $\int_{\mathbb{R}^{2}}|S|$ should be interpreted as $|S|(\mathbb{R}^{2})$ and $S^{P},S^{N},T^{P},T^{N}$ are defined by the same formula (equivalently, writing $S=\sigma|S|$ for a unit-valued $\sigma$, we have $S^{P}=(\sigma^{x}-|\sigma^{y}|)^{+}|S|$). Recall also that $\min\{\mu,\nu\}:=(\mu+\nu-|\mu-\nu|)/2$ for two real-valued measures $\mu,\nu$ (equivalently, writing $\mu=f(|\mu|+|\nu|)$ and $\nu=g(|\mu|+|\nu|)$, we have $\min\{\mu,\nu\}=\min\{f,g\}(|\mu|+|\nu|)$).
@@
$\int_{\mathbb{R}^{2}}\varphi_{r}\min\{(A_{x}^{x}-|A_{y}^{x}|)^{+},(A_{x}^{x}-|A_{y}^{x}|)^{+}\}\,d|Du|^{s}$
$\leq C\int_{\mathbb{R}^{2}}\varphi_{r}[(A_{x}^{x}-|A_{y}^{x}|)^{-}+(A_{x}^{x}-|A_{y}^{x}|)^{-}]\,d|Du|,$
@@
Given $S,T\in W^{1,1}(\mathbb{R}^{2},\mathbb{R}^{2})$, assume that $S^{x},T^{y}\geq 0$ and $\det(S,T)\geq 0$ a.e. Then
@@
The following is an immediate corollary, which will be used to prove 1.5. In its statement, we consider the class $\mathcal{M}^{+}$ of matrix-valued measures $M$ (on $\mathbb{R}^{2}$, with values in $\mathbb{R}^{2\times 2}$) such that, writing $M=A|M|$ for some unit-valued Borel $A:\mathbb{R}^{2}\to\mathbb{R}^{2\times 2}$, we have $A_{x}^{x},A_{y}^{y},\det(A)\geq 0$ ($|M|$-a.e.). We let $\sqrt{\det(M)}:=\sqrt{\det(A)}|M|$, which is a well-defined nonnegative measure.
Corollary 3.6.
Given two vector fields $S,T\in W^{1,1}(\mathbb{R}^{2},\mathbb{R}^{2})$, assume that $S^{x},T^{y}\geq 0$ and $\det(S,T)\geq 0$ a.e. Also, let $\chi:\mathbb{R}^{2}\to[0,1]$ be a Borel function supported in a bounded set. Then we have
(13) $\begin{aligned} \int_{\mathbb{R}^{2}}\chi\sqrt{\det(S,T)}&\leq\frac{1}{2}\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}\lvert\operatorname{div}S\rvert\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}\lvert\operatorname{div}T\rvert\Big]^{1/2}.\end{aligned}$
The same holds if $S^{x},S^{y},T^{x},T^{y},\operatorname{div}S,\operatorname{div}T$ are just real-valued measures on the plane with finite total variation, provided that $(S,T)\in\mathcal{M}^{+}$ (where we view $S,T$ as rows of a matrix-valued measure).
Indeed, by approximation, it is enough to check this result when $\chi\in C^{0}_{c}(\mathbb{R}^{2})$. If $S,T\in W^{1,1}(\mathbb{R}^{2},\mathbb{R}^{2})$ then the statement follows from 3.5 and Cauchy–Schwarz. In general, it follows by approximating $S,T$ by $S_{\varepsilon},T_{\varepsilon}\in C^{\infty}_{c}(\mathbb{R}^{2},\mathbb{R}^{2})$ in such a way that (viewing $S_{\varepsilon},T_{\varepsilon}$ as measures) we have the tight convergence $|S_{\varepsilon}|,|T_{\varepsilon}|\rightharpoonup|S|,|T|$, as well as $\lvert\operatorname{div}S_{\varepsilon}\rvert,\lvert\operatorname{div}T_{\varepsilon}\rvert\rightharpoonup\lvert\operatorname{div}S\rvert,\lvert\operatorname{div}T\rvert$. Indeed, by Reshetnyak’s continuity principle, we will also have
$\sqrt{\det(S_{\varepsilon},T_{\varepsilon})}\rightharpoonup\sqrt{\det(S,T)}$
as measures, since the assignment $A\mapsto\sqrt{\det(A)^{+}}$ is $1$-homogeneous.
@@
admits a $(1+C\delta)$-Lipschitz projection $\pi:B_{\delta}(\mathcal{M})\cap B_{2\Lambda}(0)\to\mathcal{M}$ for $\delta>0$ small; also, viewing $S,T$ as rows, we have $(S,T)\in\mathcal{M}$ on $B_{R+1}(0)$. Next, we can apply the standard Schoen–Uhlenbeck trick: approximating the pair $(S,T)$ by a mollification with variable radius (leaving $(S,T)$ unchanged on the complement of $B_{R+1}(0)$), by the embedding $W^{1,2}\hookrightarrow VMO$ the resulting pair $(S^{\prime},T^{\prime})$ belongs to the domain of $\pi$. We can then take $(S^{\prime\prime},T^{\prime\prime}):=\pi(S^{\prime},T^{\prime})$ as a smooth perturbation of $(S,T)$ still satisfying the assumptions.
@@
Indeed, given $\varepsilon>0$ and a nonnegative cut-off function $\varphi\in C^{\infty}_{c}(\mathbb{R}^{2})$ with $\varphi=1$ near the supports of $S$ and $T$, we can let $S^{\prime}:=\varphi\tilde{S}$ and $T^{\prime}:=\varphi\tilde{T}$, with $\tilde{S}:=S+\varepsilon\partial_{x}$ and $T^{\prime}:=T+\varepsilon\partial_{y}$. Since $S^{\prime},T^{\prime}$ are arbitrarily close to $S,T$ in the smooth topology, it is enough to prove the inequality for these new vector fields.
@@
It is easy to check that the sum $s(\mu)$ of the total variations $\lvert\operatorname{div}(\bm{1}_{\widehat{\mathcal{R}}}Z)\rvert(\mathbb{R}^{2})$, as $\widehat{\mathcal{R}}$ (and $j$) vary, equals the sum of $\int_{\partial_{L}\mathcal{R}^{\prime}}Z^{x}$ for all couples of adjacent regions $\mathcal{R},\mathcal{R}^{\prime}$ such that $\mu$ lies between $\alpha_{\mathcal{R}}$ and $\alpha_{\mathcal{R}^{\prime}}$ (regardless of the order). Hence,
@@ 3.2. A more general bound
We present here an intermediate version, which will be instrumental in obtaining 3.1.
For simplicity, assume again that $S=\alpha Z$ and $T=\beta W$, for (smooth, complete) divergence-free vector fields $Z,W$ with $Z^{x},W^{y}>0$, such that all their integral curves are graphs over the entire horizontal and vertical axes, respectively. We assume that $\alpha,\beta\in C^{\infty}_{c}(\mathbb{R}^{2})$ are nonnegative, but we drop the assumption that $\det(S,T)\geq 0$.
Given $j,k\in\mathbb{Z}$, we let $\mathcal{U}_{j}^{x}$ be the vertical stripe $j,j+1]\times\mathbb{R}$, while $\mathcal{U}_{k}^{y}$ will denote the horizontal stripe $\mathbb{R}\times[k,k+1]$.
@@
Also, given $p\in\mathcal{U}_{j}^{x}$, consider the integral curve $\gamma_{p}:\mathbb{R}\to\mathbb{R}^{2}$ of $Z$ with initial condition $\gamma_{p}(0)=p$ and let $\tau_{p}\geq 0$ be such that $\gamma_{p}^{x}(\tau_{p})=j+1$ (so that $\gamma_{p}$ leaves $\mathcal{U}_{j}^{x}$ after $\tau_{p}$). We then define $G_{j}^{x}\subseteq\mathcal{U}_{j}^{x}$ to be the (Borel) set of points $p$ such that $\gamma_{p}(t)\in p+C^{x}$ for $t\in(0,\tau_{p}]$. The set $G_{k}^{y}\subseteq\mathcal{U}_{k}^{y}$ is defined analogously (with $W$ and $C^{y}$ in place of $Z$ and $C^{x}$). Finally, we let
$G_{jk}:=G_{j}^{x}\cap G_{k}^{y}.$
The following picture illustrates a typical point $p\in G_{j}^{x}$.
@@
The key property of this set is that, whenever $p,q\in G_{jk}$, the two points cannot belong to the same integral curves for $Z$ and $W$, unless $p=q$. Indeed, if they were distinct points on the same integral curve of $Z$, then either $q\in p+C^{x}$ or $p\in q+C^{x}$; in both cases, it would follow that $|p^{y}-q^{y}|<|p^{x}-q^{x}|$. The same argument for $W$ would give the reverse inequality and thus a contradiction.
@@
holds. In order to reach this inequality, we need to localize the proof of 3.5. Bounding the integral of $\det(S_{f},T_{f})$ on the set $G_{jk}$, contained in the square $\mathcal{U}_{j}^{x}\cap\mathcal{U}_{k}^{y}$, essentially corresponds to a bound for $\int_{\mathcal{R}\cap\mathcal{S}}\det(Z,W)$ in the previous proof, for two regions $\mathcal{R}$ and $\mathcal{S}$, rather than two chains $\widehat{\mathcal{R}}$ and $\widehat{\mathcal{S}}$.
We modify the proof from the previous subsection as follows. Given any $p\in\mathcal{U}_{j}^{x}$, we can write $p=\gamma(t)$ for a unique integral curve $\gamma:[0,T]\to\mathcal{U}_{j}^{x}$ of $Z$ starting on the left side of $\mathcal{U}_{j}^{x}$ and ending on the right side (namely, $\gamma^{x}(0)=j$ and $\gamma^{x}(T)=j+1$). We then let
(17) $\alpha_{f}(p):=\min_{[0,T]}(\alpha\circ\gamma),\quad S_{f}:=\alpha_{f}Z.$
Similarly we define $\beta_{f}$ and $T_{f}$ on $\mathcal{U}_{k}^{y}$.
Since $\alpha_{f}$ is constant along integral curves of $Z$, the (continuous) vector field $\alpha_{f}Z$ is still divergence-free on $\mathcal{U}_{j}^{x}$. As in the previous subsection, we claim that
(18) $\int_{G_{jk}}\det(S_{f},T_{f})\leq\Big(\int_{\{j\}\times\mathbb{R}}S^{x}\,d\mathcal{H}^{1}\Big)\Big(\int_{\mathbb{R}\times\{k\}}T^{y}\,d\mathcal{H}^{1}\Big).$
Indeed, since $S_{f}$ is divergence-free, we can write $S_{f}=\nabla^{\perp}\varphi$ and $T_{f}=\nabla^{\perp}\psi$ for two $C^{1}$ functions $\varphi$ and $\psi$ (on $\mathcal{U}_{j}^{x}$ and $\mathcal{U}_{k}^{y}$, respectively). The integral curves of $Z$ where $\alpha_{f}>0$ correspond precisely to the regular level sets of $\varphi$.
By 3.9, we can then assert that the map $(\varphi,\psi)$ is injective on $G_{jk}\cap\{\alpha_{f}\beta_{f}>0\}$. Hence, by the area formula, $\int_{G_{jk}}\det(S_{f},T_{f})=\int_{G_{jk}}\det(\alpha_{f}Z,\beta_{f}W)$ is bounded by the area of the image of this map.
The oscillation of $\varphi$ (namely, $\sup_{\mathcal{U}_{j}^{x}}\varphi-\inf_{\mathcal{U}_{j}^{x}}\varphi$) is bounded by (and actually equal to)
$\int_{\{j\}\times\mathbb{R}}|\partial_{y}\varphi|=\int_{\{j\}\times\mathbb{R}}\alpha_{f}Z^{x}\leq\int_{\{j\}\times\mathbb{R}}\alpha Z^{x}=\int_{\{j\}\times\mathbb{R}}S^{x},$
thanks to the fact that $\alpha_{f}\leq\alpha$. This, together with the same bound for the oscillation of $\psi$, gives our claim Eq. 18.
Further, note that
$\int_{\{j\}\times\mathbb{R}}S^{x}=\int_{\{s\}\times\mathbb{R}}S^{x}-\int_{[j,s]\times\mathbb{R}}\operatorname{div}S$
for all $s\in[j,j+1]$. Averaging over this interval, we obtain

### v2

@@ Abstract.
While a monotonicity formula is not available for critical points [4], when $k=2$ and $n=3$ we show that the Michael–Simon inequality holds if $F$ is close to $1$ (in $C^{1}$), meaning that $\mathcal{F}$ is close to the usual area.
@@ 1.1. Setting and main result
In particular, if $\theta\geq\theta_{0}$ $|V|$-a.e., Eq. 1 holds for a possibly different constant $C(F)$.
@@ 1.2. New variants of multilinear Kakeya in dimension two
Given $S,T\in W^{1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, assume that $S^{x},T^{y}\geq 0$ and $\det(S,T)\geq 0$ a.e. Then
@@
Also, let $\chi:\mathbb{R}^{2}\to[0,1]$ be a Borel function supported in a bounded set. Then, for some universal constants $C,C^{\prime}>0$, we have
(4) $\begin{aligned} \int_{\mathbb{R}^{2}}\chi\min\{S^{P},T^{P}\}&\leq C\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}(|S|+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(|T|+\lvert\operatorname{div}T\rvert)\Big]^{1/2}\\ &\quad+C\int_{\mathbb{R}^{2}}(\lvert\operatorname{div}S\rvert+\lvert\operatorname{div}T\rvert)+C^{\prime}\int_{\mathbb{R}^{2}}(S^{N}+T^{N}).\end{aligned}$
The same holds if $S^{x},S^{y},T^{x},T^{y},\operatorname{div}S,\operatorname{div}T$ are just real-valued measures on the plane with finite total variation. Moreover, if $S^{x},T^{y}\geq 0$ then we can take $C^{\prime}=1$.
@@ 1.3. General results in arbitrary dimension
While we study autonomous integrands $F(P)$, for simplicity and in order to have scale-invariant bounds when possible, in general one can consider non-autonomous ones, of the form $F(x,P)$. This generalization is necessary in order to look at anisotropic integrands on manifolds $M^{n}$, where $F$ is defined on the Grassmannian bundle of $k$-planes and thus locally takes this form. Compared to the autonomous case, the general case just introduces an error term which is easily handled (letting ${\tilde{\delta}}^{F}V$ denote the right-hand side of Eq. 6, we have $|{\tilde{\delta}}^{F}V|\leq|\delta^{F}V|+C|V|$ locally). Straightforward modifications show that all the stated consequences of Michael–Simon still hold true, assuming this bound in a local form such as
@@ Acknowledgements
As already mentioned, this paper uses many ideas due to Frederick J. Almgren Jr. The authors are very grateful to Jean Taylor, who shared with them the unpublished manuscript of Almgren. They also wish to thank William Cooperman and Adolfo Arroyo-Rabasa for identifying an issue in an earlier draft, which prompted a more general formulation of 1.7. A.P.’s research is funded by the European Research Council (ERC) through StG 101165368 “MAGNETIC”. G.D.P.’s research is funded by the European Research Council (ERC) through CoG 101169953 “RISE”. [^1] Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council.
@@
Up to a subsequence, we can assume that $|\delta^{F}V_{i}|\rightharpoonup\nu$ for a suitable Radon measure $\nu$. For $|V|$-a.e. $x$, the rectifiable varifold $V$ admits an approximate tangent plane and there exists $C>0$, depending on $x$, such that [^2] Recall that, for any two Radon measures $\nu,\mu\geq 0$, we have $\limsup_{r\to 0}\frac{\nu(B_{r}(x))}{\mu(B_{r}(x))}<\infty\quad\text{for $\mu$-a.e.\ $x$.}$
@@ 3. Nonlinear inequalities for vector fields on the plane
Also, let $\chi:\mathbb{R}^{2}\to[0,1]$ be a Borel function supported in a bounded set. Then, for some universal constants $C,C^{\prime}>0$, we have
(12) $\begin{aligned} \int_{\mathbb{R}^{2}}\chi\min\{S^{P},T^{P}\}&\leq C\|\chi\|_{L^{2}}\Big[\int_{\mathbb{R}^{2}}(|S|+\lvert\operatorname{div}S\rvert)\Big]^{1/2}\Big[\int_{\mathbb{R}^{2}}(|T|+\lvert\operatorname{div}T\rvert)\Big]^{1/2}\\ &\quad+C\int_{\mathbb{R}^{2}}(\lvert\operatorname{div}S\rvert+\lvert\operatorname{div}T\rvert)+C^{\prime}\int_{\mathbb{R}^{2}}(S^{N}+T^{N}).\end{aligned}$
The same holds if $S^{x},S^{y},T^{x},T^{y},\operatorname{div}S,\operatorname{div}T$ are just real-valued measures on the plane with finite total variation. Moreover, if $S^{x},T^{y}\geq 0$ then we can take $C^{\prime}=1$.
@@
In the general case of measures, quantities such as $\int_{\mathbb{R}^{2}}|S|$ should be interpreted as $|S|(\mathbb{R}^{2})$ and $S^{P},S^{N},T^{P},T^{N}$ are defined by the same formulas Eq. 11 (equivalently, writing $S=\sigma|S|$ for a unit-valued $\sigma$, we have $S^{P}=(\sigma^{x}-|\sigma^{y}|)^{+}|S|$). Recall also that $\min\{\mu,\nu\}:=(\mu+\nu-|\mu-\nu|)/2$ for two real-valued measures $\mu,\nu$ (equivalently, writing $\mu=f(|\mu|+|\nu|)$ and $\nu=g(|\mu|+|\nu|)$, we have $\min\{\mu,\nu\}=\min\{f,g\}(|\mu|+|\nu|)$).
@@
$\int_{\mathbb{R}^{2}}\varphi_{r}\min\{(A_{y}^{y}-|A_{x}^{y}|)^{+},(A_{x}^{x}-|A_{y}^{x}|)^{+}\}\,d|Du|^{s}$
$\leq C\int_{\mathbb{R}^{2}}\varphi_{r}[(A_{y}^{y}-|A_{x}^{y}|)^{-}+(A_{x}^{x}-|A_{y}^{x}|)^{-}]\,d|Du|,$
@@
Given $S,T\in W^{1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, assume that $S^{x},T^{y}\geq 0$ and $\det(S,T)\geq 0$ a.e. Then
@@
admits a $(1+C\delta)$-Lipschitz projection $\pi:B_{\delta}(\mathcal{M})\cap B_{2\Lambda}(0)\to\mathcal{M}$ for $\delta>0$ small; also, viewing $S,T$ as rows, we have $(S,T)\in\mathcal{M}$ on $B_{R+1}(0)$.
Next, we can apply the standard Schoen–Uhlenbeck trick: approximating the pair $(S,T)$ by a mollification with variable radius (leaving $(S,T)$ unchanged on the complement of $B_{R+1}(0)$), by the embedding $W^{1,2}\hookrightarrow VMO$ the resulting pair $(S^{\prime},T^{\prime})$ belongs to the domain of $\pi$ (on $B_{R+1}(0)$). We can then take $(S^{\prime\prime},T^{\prime\prime}):=\pi(S^{\prime},T^{\prime})$ as a smooth perturbation of $(S,T)$ (letting it equal to $(S^{\prime},T^{\prime})$ outside of $B_{R+1}(0)$) still satisfying the assumptions: since
$|(S^{\prime\prime},T^{\prime\prime})-(S^{\prime},T^{\prime})|\leq C\delta,\quad|d(S^{\prime\prime},T^{\prime\prime})|^{2}\leq(1+C\delta)|d(S^{\prime},T^{\prime})|^{2},$
the pair $(S^{\prime\prime},T^{\prime\prime})$ converges strongly in $W^{1,2}(\mathbb{R}^{2})$ to $(S^{\prime},T^{\prime})$, as $\delta\to 0$.
@@
Indeed, given $\varepsilon>0$ and a nonnegative cut-off function $\varphi\in C^{\infty}_{c}(\mathbb{R}^{2})$ with $\varphi=1$ near the supports of $S$ and $T$, we can let $S^{\prime}:=\varphi\tilde{S}$ and $T^{\prime}:=\varphi\tilde{T}$, with $\tilde{S}:=S+\varepsilon\partial_{x}$ and $\tilde{T}:=T+\varepsilon\partial_{y}$. Since $S^{\prime},T^{\prime}$ are arbitrarily close to $S,T$ in the smooth topology, it is enough to prove the inequality for these new vector fields.
@@
It is easy to check that the sum $s(\mu)$ of the total variations $\lvert\operatorname{div}(\bm{1}_{\widehat{\mathcal{R}}}Z)\rvert(\mathbb{R}^{2})$, as $\widehat{\mathcal{R}}$ (and $j$) vary, equals the sum of $\int_{\partial_{L}\mathcal{R}^{\prime}}Z^{x}$ for all pairs of adjacent regions $\mathcal{R},\mathcal{R}^{\prime}$ such that $\mu$ lies between $\alpha_{\mathcal{R}}$ and $\alpha_{\mathcal{R}^{\prime}}$ (regardless of the order). Hence,
@@ 3.2. A more general bound
We present here an intermediate version, which will be instrumental in obtaining 3.1. Let us start by introducing some additional terminology. Given $j,k\in\mathbb{Z}$, we let
$\mathcal{U}_{j}^{x}:=[j,j+1]\times\mathbb{R},\quad\mathcal{U}_{k}^{y}:=\mathbb{R}\times[k,k+1],$
called vertical and horizontal stripes, respectively. Assume now that $S,T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ are smooth vector fields.
Definition 3.7.
We say that $p\in\{j\}\times\mathbb{R}$ is a good initial condition for $S$ if, for the integral curve $\gamma_{p}$ of $S$ with $\gamma_{p}(0)=p$, there exists $\tau_{p}>0$ such that
$\gamma_{p}(\tau_{p})\in\{j+1\}\times\mathbb{R},\quad\gamma_{p}(t)\in(j,j+1)\times\mathbb{R}\text{ for all }t\in(0,\tau_{p}),$
as well as $\dot{\gamma}_{p}^{x}(0),\dot{\gamma}_{p}^{x}(\tau_{p})>0$. We say that $\mathcal{R}^{x}\subset\mathcal{U}_{j}^{x}$ is a good region for $S$ if there exists a segment $\mathcal{I}:=\{j\}\times[a,b]$ (with $a<b$) on the left side, consisting of good initial conditions for $S$, such that
$\mathcal{R}^{x}=\bigcup_{p\in\mathcal{I}}\gamma_{p}([0,\tau_{p}]).$
The same notions for $T$ are defined analogously, interchanging the roles of $x,y$.
Note that, since each $\gamma_{p}$ meets $\partial\mathcal{U}_{j}^{x}$ transversely, $\tau_{p}$ depends smoothly on $p$. In particular, a good region $\mathcal{R}^{x}$ is diffeomorphic to a compact rectangle $a,b]\times[0,1]$ and $S(q)\neq 0$ for all $q\in\mathcal{R}^{x}$.
Given two good regions $\mathcal{R}^{x},\mathcal{R}^{y}$ for $S,T$, respectively, as in the proof of 3.5 we can construct smooth vector fields $Z:\mathcal{R}^{x}\to\mathbb{R}^{2}$ and $W:\mathcal{R}^{y}\to\mathbb{R}^{2}$, as well as smooth functions $\alpha:\mathcal{R}^{x}\to\mathbb{R}$ and $\beta:\mathcal{R}^{y}\to\mathbb{R}$, such that
$\operatorname{div}Z=0,\quad\operatorname{div}W=0,\quad S=\alpha Z,\quad T=\beta W,\quad\alpha>0,\quad\beta>0.$
Note that we are no longer assuming that $S^{x}\geq 0$ and $T^{y}\geq 0$, nor that $\det(S,T)\geq 0$ on $\mathcal{R}^{x}\cap\mathcal{R}^{y}$ (although we do have $S^{x}>0$ on the left and right sides of $\mathcal{R}^{x}$, and similarly for $W$).
@@
Also, given $p\in\mathcal{R}^{x}$, consider the (maximal) integral curve $\gamma_{p}:I\to\mathbb{R}^{2}$ of $S$ with initial condition $\gamma_{p}(0)=p$ and let $\tau_{p}\geq 0$ be the first time such that $\gamma_{p}^{x}(\tau_{p})=j+1$. We then define $G^{x}\subseteq\mathcal{R}^{x}$ to be the Borel set of points $p$ such that $\gamma_{p}(t)\in p+C^{x}$ for all $t\in(0,\tau_{p}]$. The set $G^{y}\subseteq\mathcal{R}^{y}$ is defined analogously (with $T$ and $C^{y}$ in place of $S$ and $C^{x}$). Finally, we let
$G:=G^{x}\cap G^{y}.$
The following picture illustrates a typical point $p\in G^{x}$.
@@
The key property of this set is that, whenever $p,q\in G$, the two points cannot belong to the same integral curves for $S$ and $T$, unless $p=q$. Indeed, if they were distinct points on the same integral curve of $S$, then either $q\in p+C^{x}$ or $p\in q+C^{x}$; in both cases, it would follow that $|p^{y}-q^{y}|<|p^{x}-q^{x}|$. The same argument for $T$ would give the reverse inequality and thus a contradiction.
@@
holds. In order to reach this inequality, we need to localize the proof of 3.5, as follows.
Given any $p\in\mathcal{R}^{x}$, we can write $p=\gamma(t)$ for a unique (piece of) integral curve $\gamma:[0,T]\to\mathcal{R}^{x}$ of $Z$ starting on the left side $\partial_{L}\mathcal{R}^{x}$ of $\mathcal{R}^{x}$ and ending on the right side (namely, $\gamma^{x}(0)=j$ and $\gamma^{x}(T)=j+1$). We then let
(16) $\alpha_{f}(p):=\min_{[0,T]}(\alpha\circ\gamma)>0,\quad S_{f}:=\alpha_{f}Z.$
Similarly we define $\beta_{f}$ and $T_{f}$ on $\mathcal{R}^{y}$.
Since $\alpha_{f}$ is constant along integral curves of $Z$, the (continuous) vector field $\alpha_{f}Z$ is still divergence-free on $\mathcal{R}^{x}$. As in the previous subsection, we claim that
(17) $\int_{G}\det(S_{f},T_{f})\leq\Big(\int_{\partial_{L}\mathcal{R}^{x}}S^{x}\,d\mathcal{H}^{1}\Big)\Big(\int_{\partial_{B}\mathcal{R}^{y}}T^{y}\,d\mathcal{H}^{1}\Big),$
where the last two integrals are on the left and bottom sides, respectively. Indeed, since $S_{f}$ is divergence-free and $\mathcal{R}^{x}$ is simply connected, we can write $S_{f}=\nabla^{\perp}\varphi$, and similarly $T_{f}=\nabla^{\perp}\psi$, for two $C^{1}$ functions $\varphi$ and $\psi$ (on $\mathcal{R}^{x}$ and $\mathcal{R}^{y}$, respectively).
Since $S_{f}^{x}>0$ on the left side $\partial_{L}\mathcal{R}^{x}$, $\varphi$ is injective here, and hence each integral curve of $S$ (in $\mathcal{R}^{x}$) is an entire level set of $\varphi$; the same holds for $T$ and $\psi$. By 3.9, we can then assert that the map $(\varphi,\psi)$ is injective on $G$. Hence, by the area formula, $\int_{G}\det(S_{f},T_{f})=\int_{G}\det(\alpha_{f}Z,\beta_{f}W)$ is bounded by the area of the image of this map.
The oscillation of $\varphi$ (namely, $\sup_{\mathcal{R}^{x}}\varphi-\inf_{\mathcal{R}^{x}}\varphi$) is bounded by (and actually equal to)
$\int_{\partial_{L}\mathcal{R}^{x}}|\partial_{y}\varphi|=\int_{\partial_{L}\mathcal{R}^{x}}\alpha_{f}Z^{x}\leq\int_{\partial_{L}\mathcal{R}^{x}}\alpha Z^{x}=\int_{\partial_{L}\mathcal{R}^{x}}S^{x},$
where we used the fact that $\alpha_{f}\leq\alpha$. This, together with the same bound for the oscillation of $\psi$, gives our claim Eq. 17.
Further, since $S$ is tangent to the top and bottom sides of $\mathcal{R}^{x}$ (namely, to the boundary $\partial\mathcal{R}^{x}$ in the open stripe $(j,j+1)\times\mathbb{R}$), we have
$\int_{(\{j\}\times\mathbb{R})\cap\mathcal{R}^{x}}S^{x}=\int_{(\{s\}\times\mathbb{R})\cap\mathcal{R}^{x}}S^{x}-\int_{([j,s]\times\mathbb{R})\cap\mathcal{R}^{x}}\operatorname{div}(S)$
for a.e. $s\in[j,j+1]$. Averaging over this interval, we obtain